Employees of a discount appliance store receive an additional 20% off of the lowest price on an item. If an employee purchases a dishwasher during a 15% off sale, how much will he pay if the dishwasher originally cost $450?how do they get 306$
Answer by dabanfield(803)

Employees of a discount appliance store receive an additional 20% off of the lowest price on an item. If an employee purchases a dishwasher during a 15% off sale, how much will he pay if the dishwasher originally cost $450?how do they get 306$

If the original amount is $450. The first discount is
.15 * 450 = 67.50
The amount after the first discount is then 450 - 67.50 = 382.50
So we apply the second discount (i.e. 20%) to the 382.50:
.20 * 382.50 = 76.50
The fully discounted amount is then 382.50 - 76.50 = $306